# Error thresholds of toric codes with transversal logical gates

### v1

@@ II.1 Stat-mech mapping with perfect syndromes
where $C$ denotes a “cycle” of Pauli operators that do not trigger any syndromes when the stabilizers are measured [^2] We call $C$ a cycle since in 2D topological codes such an operator is usually supported on closed cycles.. That is, it either belongs to the stabilizer group of the code state, denoted by $S$, or the set of non-trivial logical operator, denoted by $L$. In this way, Pauli errors $E$ and $E\cdot C$ will trigger exactly the same set of syndromes. In the first equality, the expression in square brackets is the probability of logical error given that a physical error $E$ occurs, which is computed via the second equality. Without the knowledge of the ground truth about $E$, the optimal maximum likelihood decoder guesses an error $E\cdot C$ according to the conditional probability $\mathrm{prob}(E\cdot C|E)$. A logical error will occur whenever our guess differs from the actual physical error by any logical operator, see Fig. 1(a). Averaging over all possible errors yields the logical error rate in Eq. (II.1).
@@ III.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
The stat-mech model allows us to form a theoretical anticipation of the error thresholds of the two code blocks, even before carrying out the numerical simulation. In light of Eq. (4), the logical error rate for the control/target blocks now maps to the sum over $\overline{e^{-F_{DW}^{\sigma/\tau}}}$, where $F^{\sigma/\tau}_{DW}$ is the domain wall free energy of a global domain wall of the spins $\sigma/\tau$, and the disorder average is taken over both $E^{c}$ and $E^{t}$ (i.e. $s^{c}$ and $s^{t}$). Therefore, the error thresholds of the control and target code blocks, $p^{c}_{\text{c}}$ and $p^{t}_{\text{c}}$, correspond to the order-disorder phase transition of the $\sigma$ and $\tau$ spins, respectively. From Eq. (18), we expect that the target block will first reach the error threshold with increasing $\tilde{p}$. Since the critical point is within the ordered phase of $\{\sigma\}$ spins, we expect the coupling term between $\{\sigma\}$ and $\{\tau\}$ spins to enhance the error threshold of the target block $p^{t}_{\text{c}}$. On the other hand, since $K_{2}<J$, the error threshold of the control layer $p^{c}_{\text{c}}$ should also be reduced compared to the case of quantum memory, $p_{\text{th}}\approx 0.109$.
@@ V Conclusion and outlook
Figure 10: (a) A noisy transversal Clifford logical circuit between two code blocks $A$ and $B$ consisting of two tCNOT gates and a (fold-)transversal $S$ gate. Ever pair of transversal gates are separated by $\Delta t$ rounds of syndrome error extractions and generic physical error channels $\mathcal{N}$. (b) Sketch of the impact of transversal gates on the stat-mech model that is mapped from the circuit in (a). The $X/Z$ spacetime detectors of the two code blocks $A$ and $B$, denoted by $\{d^{A/B}_{X/Z}\}$, give rise to 4 different set of classical spins in total, which are coupled at the plane defects induced by the logical operations. The propagation of Pauli errors by the transversal gates are marked by red arrows. The impact region $|t-T|<\lambda$ due to increased effective local temperature are marked by darker colors for each plot.
@@
Throughout this work, we specifically consider logical circuits with a single tCNOT gate. However, our study can be straightforwardly generalized to map a general transversal Clifford logical circuit of toric codes to a stat-mech model, where each consecutive logical operation is separated by $\Delta t$ rounds of syndrome extractions, see Fig. 10(a). In this stat-mech model of the logical circuit, each transversal Clifford gate operation affects a specific time slice in the 3D stat-mech model. Hence, the effect of each gate at $t=T$ will be local in the time direction inside a time window $|t-T|\leq\lambda$, where $\lambda\sim O(1)$, as shown pictorially in Fig. 10(b). For a given circuit, one should be able to simulate the corresponding models based on the mappings that we have provided in this paper. Therefore, our estimation of error threshold can be readily generalized to estimate the threshold of the entire logical circuit.
@@ Appendix A Coupling constants in the Hamiltonian of the random Ashkin-Teller model
Appendix B Derivation of the stat-mech model with syndrome errors
@@
$r^{c}_{p}(t)\to r^{c}_{p}\prod_{l\in p}\sigma_{l}(t),\ r^{t}_{p}(t)\to r^{t}_{p}\prod_{l\in p}\tau_{l}(t),$ (31)

### v2

@@ I Introduction
CSS code + transversal Clifford Generic Pauli Yes See Appendix D
@@ II.1 Stat-mech mapping with perfect syndromes
where $C$ denotes a “cycle” of Pauli operators that do not trigger any syndromes when the stabilizers are measured [^2] We call $C$ a cycle since in 2D topological codes such an operator is usually supported on closed cycles.. That is, it either belongs to the stabilizer group of the code state, denoted by $S$, or the set of non-trivial logical operator, denoted by $L$. In this way, Pauli errors $E$ and $E\cdot C$ will trigger exactly the same set of syndromes. In the first equality, the expression in square brackets is the probability of logical error given that a physical error $E$ occurs, which is computed via the second equality. Without the knowledge of the ground truth about $E$, the optimal decoder guesses an error $E\cdot C$ according to the conditional probability $\mathrm{prob}(E\cdot C|E)$. A logical error will occur whenever our guess differs from the actual physical error by any logical operator, see Fig. 1(a). Averaging over all possible errors yields the logical error rate in Eq. (II.1).
@@ III.3 Stat-mech mapping for perfect syndromes and persistent bit-flip errors
The stat-mech model allows us to form a theoretical anticipation of the error thresholds of the two code blocks, even before carrying out the numerical simulation. In light of Eq. (4), the logical error rate for the control/target blocks now maps to the sum over $\overline{e^{-F_{DW}^{\sigma/\tau}}}$, where $F^{\sigma/\tau}_{DW}$ is the domain wall free energy of a global domain wall of the spins $\sigma/\tau$, and the disorder average is taken over both $E^{c}$ and $E^{t}$ (i.e. $s^{c}$ and $s^{t}$). Therefore, the error thresholds of the control and target code blocks, $p^{c}_{\text{c}}$ and $p^{t}_{\text{c}}$, correspond to the order-disorder phase transition of the $\sigma$ and $\tau$ spins, respectively. From Eq. (18), we expect that the target block will first reach the error threshold with increasing $\tilde{p}$. Since the critical point is within the ordered phase of $\{\sigma\}$ spins, we expect the coupling term between the $\{\sigma\}$ and $\{\tau\}$ spins to enhance the error threshold of the target block $p^{t}_{\text{c}}$. On the other hand, since $K_{2}<J$, the error threshold of the control layer $p^{c}_{\text{c}}$ should also be reduced compared to the case of quantum memory, $p_{\text{th}}\approx 0.109$.
@@ V Conclusion and outlook
Figure 10: (a) A noisy transversal Clifford logical circuit between two code blocks $A$ and $B$ consisting of two tCNOT gates and a (fold-)transversal $S$ gate. Ever pair of transversal gates are separated by $\Delta t$ rounds of syndrome error extractions and generic physical error channels $\mathcal{N}$. (b) Sketch of the impact of transversal gates on the stat-mech model that is mapped from the circuit in (a). The $X/Z$ spacetime detectors of the two code blocks $A$ and $B$, denoted by $\{d^{A/B}_{X/Z}\}$, give rise to 4 different set of classical spins in total (see Appendix C for details), which are coupled at the plane defects induced by the logical operations. The propagation of Pauli errors by the transversal gates are marked by red arrows. The impact region $|t-T|<\lambda$ due to increased effective local temperature are marked by darker colors for each plot.
@@
Throughout this work, we specifically consider toric code logical circuits with a single tCNOT gate. However, our study can be straightforwardly generalized to map a general transversal Clifford logical circuit of toric codes to a stat-mech model, where each consecutive logical operation is separated by $\Delta t$ rounds of syndrome extractions, see Fig. 10(a). In Appendix C, we carry out detailed stat-mech mapping of fold-transversal Hadamard and $S$ gate for toric code. Combined with the tCNOT gate discussed in the main text, we are able to derive stat-mech model of the full transversal Clifford logical circuit. Crucially, each transversal Clifford gate operation affects a specific time slice in the 3D stat-mech model. Hence, the effect of each gate at $t=T$ will be local in the time direction inside a time window $|t-T|\leq\lambda$, where $\lambda\sim O(1)$, as shown pictorially in Fig. 10(b). For a given circuit, one should be able to simulate the corresponding models based on the mappings that we have provided in this paper. Therefore, our estimation of error threshold can be readily generalized to estimate the threshold of the entire logical circuit.
Our stat-mech approach can even be generalized to arbitrary Calderbank-Steane-Shor (CSS) codes that possess transversal logical gates. We refer interested readers to Appendix D for more details.
@@ Appendix A Coupling constants in the Hamiltonian of the random Ashkin-Teller model
Appendix B Derivation of the stat-mech model for tCNOT with syndrome errors
@@
$r^{c}_{p}(t)\to r^{c}_{p}(t)\prod_{l\in p}\sigma_{l}(t),\ r^{t}_{p}(t)\to r^{t}_{p}(t)\prod_{l\in p}\tau_{l}(t),$ (31)